# Joint Identity Verification and Pose Alignment for Partial Fingerprints

### v2

@@ Abstract
Currently, portable electronic devices are becoming more and more popular. For lightweight considerations, their fingerprint recognition modules usually use limited-size sensors. However, partial fingerprints have few matchable features, especially when there are differences in finger pressing posture or image quality, which makes partial fingerprint verification challenging. Most existing methods regard fingerprint position rectification and identity verification as independent tasks, ignoring the coupling relationship between them — relative pose estimation typically relies on paired features as anchors, and authentication accuracy tends to improve with more precise pose alignment. Consequently, in this paper we propose a method for joint identity verification and pose alignment of partial fingerprint pairs, aiming to leverage their inherent correlation to improve each other. To achieve this, we propose a multi-task CNN (Convolutional Neural Network)-Transformer hybrid network, and design a pre-training task to enhance the feature extraction capability. Experiments on multiple public datasets (NIST SD14, FVC2002 DB1_A & DB3_A, FVC2004 DB1_A & DB2_A, FVC2006 DB1_A) and an in-house dataset demonstrate that our method achieves state-of-the-art performance in both partial fingerprint verification and relative pose estimation, while being more efficient than previous methods.
@@ I Introduction
Figure 1: Partial fingerprint matching faces three challenges: sparse minutiae distribution, similar local texture and significant modal difference. This figure shows corresponding representative examples: (a) Three weakly-textured fingerprints that lack appropriate landmark feature points. (b) Local ridge patterns of fingerprints from different fingers have high similarity. (c) Fingerprint features in different modalities (in this case from left to right are dry, normal and wet respectively) may be missed (green) or incorrectly extracted (red). Visualized feature points are minutiae extracted by VeriFinger [9].
A natural idea for partial fingerprint matching is to transfer existing solutions for rolled or plain fingerprint recognition. Nevertheless, conventional algorithms usually rely on a sufficient number of key points to establish and compare spatial structural relationships [10, 11]. While certain studies use neural networks to extract fixed-length region descriptors [12, 13, 14, 15], consequently reducing the dependence on point features, these approaches require either large effective area or exceedingly precise alignment. A series of fusion strategies are subsequently proposed at feature [12, 16], image [17, 18] and decision levels [19, 20], but most of them still depend on one-to-one matching or require several densely distributed high-quality images. In summary, although these highly generalized feature forms ensure efficient matching, some local details are missed, which is not enough to support high-precision recognition for partial fingerprints due to their extremely small effective area. Considering this, some researchers simultaneously input fingerprint images and extract interrelated features to fully utilize complete information [7, 21], which effectively improves the matching performance on partial fingerprints. However, to the best of our knowledge, almost all existing fingerprint matching methods (compared in Fig. 2) regard pose rectification and identity verification as independent tasks, ignoring the coupling relationship between them. Specifically, the estimation of relative pose typically relies on paired regional features as anchors, and the accuracy of authentication generally increases with enhanced precision in pose alignment. This is similar to a pictorial jigsaw puzzle where joint decision is suggested about whether (are they neighboring image fragments? ) and how (what’s the reassemble solution? ) any two pieces fit together [22, 23], in order to better balance robustness and speed in the overall process, rather than being made individually.
@@ II Related Work
According to different execution frameworks and feature representation forms, existing one-to-one fingerprint matching algorithms can be mainly divided into three categories: key points based, fixed-length region descriptor based, and pairwise interrelated feature based. In the first two types of algorithms highly generalized features are stored and compared here (in other words, each fingerprint undergoes feature extraction only once) to enable efficient searches in large-scaled galleries. On the other hand, pairwise interrelated feature based methods could use richer information (features need to be re-extracted for each match) and is more suitable for confirmation scenarios with high precision requirements and few gallery / reference fingerprints, such as unlocking a mobile phone. Given the multitude of fingerprint matching algorithms, this rough classification may not comprehensively encompass all works. Nonetheless, we have endeavored to summarize them to the best of our knowledge. Fig. 2 shows the corresponding schematic diagrams of these processes.
@@ II-A Key Points Based
Minutiae are the most popular features used for fingerprint matching [3]. Researchers define minutiae descriptors using the attributes of minutiae themselves [10] and auxiliary information (such as orientation [32], period map [33], ridge [11] etc.). Considering the lack of reliable minutiae in latent fingerprints, Cao et al.[19] uniformly sampled images into small patches and used a network to extract fixed length vector to describe virtual minutiae. However, as explained in Fig. 1, these features are not discriminative enough in partial fingerprints. On the other hand, some studies introduced universal key points in computer vision field into fingerprints [34, 6]. Features extracted in this way are more densely distributed, but at the same time more sensitive to unreliable image details and therefore easily confused. There are also algorithms that exploit Level-3 features from high-resolution images [35, 36], which are relatively expensive to deploy in most commercial applications. After getting key point sets, correspondence between them will be established based on the similarity of features and spatial structures, and then relative poses and matching scores are calculated.
@@ II-B Fixed-length Region Representation Based
Early studies use hand-crafted statistics to characterize fingerprint regions and convert them into fixed length vectors [37, 38, 39]. This feature form significantly improves the speed of large database indexing at the expense of reduced accuracy, and is usually used for rough searching. Engelsma et al.[12] propose an end-to-end network to extract fixed-length global representations of minutiae and textures, which greatly improves the performance of fixed-length representation methods. Inspired by this, researchers have conducted more exploration and improvement of fixed-length representation methods based on deep learning [14, 40, 15]. In addition, Gu et al.[13] conducted dense sampling on complete fingerprints and compared the fixed length descriptors of each patch, ultimately making a comprehensive decision. These algorithms usually require pose rectification in advance, such as adaptively adjusting images through a spatial transformation layer [12, 15], rectifying absolute pose of fingerprints through a separate network [40], or exploiting the positioning relationships implicit in minutiae [14]. The error of pose estimation is tolerated to a certain extent because the extracted features are a generalization of regional characteristics. However, current methods may struggle to address partial fingerprints, as the pose estimation on them is often unstable.
@@ II-C Pairwise Interrelated Feature Based
Compared with key points or fixed-length descriptors, gray-level features of pixels contain richer and more discriminatory information about local textures [41]. Traditional algorithms use image correlation for fingerprint verification, which can be affected by skin distortion [41, 42]. Some works on fingerprint distortion further increase the tolerance of ridge misalignment [43]. With the development of deep learning, researchers apply neural networks to directly extract and verify interrelated features between each other, significantly improving the verification performance and efficiency [44, 45, 46]. He et al.[7] use multi-rotation and multi-size cropped image pairs as combined inputs to help the network better understand scenarios with large spatial transformations or small overlapping areas. It should be noted that a significant difference between this type of method and methods based on fixed-length region descriptor is that the relative pose of a certain fingerprint is rectified instead of the respective absolute pose. Features of such methods are then compared more precisely with spatial correspondence thus have stricter requirements for alignment. Obviously, the verification performance will significantly decline with inaccurate pose estimation [7].
@@ III Method
In this paper, we propose JIPNet, a CNN-Transformer hybrid network, to jointly estimates the authentication probability and relative pose of partial fingerprint pairs. Fig. 3 gives the complete flowchart of our proposed algorithm. The overall process can be divided into four phases, which are indicated by different color bars in schematic: independent feature extraction (red), interrelated feature extraction (gold), identity verification (blue) and pose prediction (green). For an input fingerprint pair, our algorithm first extracts the respective features in parallel using convolutional blocks with shared weights (in Stem and Stages 1, 2). These two sets of features are then concatenated and fed to transformer blocks in Stage 3 (along with some extra microprocessing) for sufficient information interaction both locally and globally. Finally, respective output layers utilize the aggregated information to predict classification probabilities for authentication (Stage 4 - C and Classifier) and regression values for relative poses (Stage 4 - R and Regressor). It should be noted that these two tasks are estimated in an integrated manner instead of being performed independently as previous methods because we hope to utilize the mutually beneficial coupling relationship between each other, inspired by works on jigsaw puzzles [22, 23]. In particular, pose-related regression layers could be pruned to improve efficiency since only identity identification is necessary in most applications.
@@ III-B Interrelated Feature Extraction
The hybrid strategy of using early CNN and subsequent transformer to aggregate local details and global information has achieved great success in multiple fields [24, 25, 26]. Inspired by these, we utilize transformers to capture long-range dependencies within and across paired features, thereby ensuring sufficient information interaction. The paired features $f_{1}$ and $f_{2}$ are first biased with a sinusoidal positional encoding term [24], and then spatially flattened into 1D sequences and fed into stacked transformers with interleaved self- and cross- attention. Let $x$ and $y$ denote the 2D position, $i$ and $d$ denote the current and total dimensions, the additional absolute position information is calculated as:
@@
$PE_{(x,y,4i+3)}=\cos\left(y/10000^{4i/d}\right)\;.$
It should be noted that the next two attention blocks have the same structure, including residual-connected Multi-Head Attention (MHA) and Feed-Forward Network (FFN) [50], only the data streams are differentiated by their respective expected effects. Specifically, let $\operatorname{SA}\left(f_{i},f_{j}\right)$ (Self-Attention) and $\operatorname{CA}\left(f_{i},f_{j}\right)$ (Cross-Attention) represent the corresponding sub-blocks in Fig. 3, the complete process of transformer block is defined as:
@@ III-C Identification Verification & Pose Prediction
Similar to the previous phase, features are separately fed into stacked identical transformer blocks with head of $12$ to prepare for the corresponding task head. We define identity verification as a binary classification task, that is, judging whether the input image pairs come from a same finger. Let $\mathcal{F}$, $\operatorname{BN}$, $\operatorname{Swish}$ and $\operatorname{AvgPool}$ denote fully connected layer, batch normalization, swish activation and adaptive average pooling respectively, the classification probability $p$ is calculated as:
@@ V-E Visual Analysis
In order to provide more specific and intuitive comparison with interpretability, we employ occlusion sensitivity [21] to approximate the contribution of distinct local areas to the overall matching judgment. Fig. 10 shows four representative visualization results. Query fingerprint in images pairs is centered and search fingerprint is rigidly aligned based on the ground truth as background. Subsequently, a $32\times 32$ mask is slid and applied to search fingerprints and the differences between scores before and after occlusion are calculated, which are normalized and overlaid as the heatmap of search fingerprint. Considering that there is no reference relationship between the matching scores directly obtained (or mapped) by different algorithms, we convert these scores into a comparable probability form and attached them in corresponding visualization results. Specifically, given the genuine matching score $s$ and overall distributions (shown in Fig. 7), the verification accuracy can be expressed as:
$P\left(y=1|s\right)$ (15)
$=\frac{P\left(s|y=1\right)P\left(y=1\right)}{P\left(s|y=1\right)P\left(y=1\right)+P\left(s|y=0\right)P\left(y=0\right)}$
$=\frac{f_{\mathrm{1}}\left(s\right)P\left(y=1\right)}{f_{\mathrm{1}}\left(s\right)P\left(y=1\right)+f_{\mathrm{0}}\left(s\right)P\left(y=0\right)}\;,$
where $P$ and $f$ represent the probability and probability density, $y$ represents the identity to be verified (1 is the same, 0 is different). It can be seen that algorithms based on key points [6, 9] and region representation [12, 13] expose their respective shortcomings when there is a lack of sufficient features in overlapping areas (row 1) and significant differences in ridge texture of non overlapping areas (row 2). PFVNet [7] has clear advantages in both cases, even with minimal overlap (row 3). However, it fails in the last example with confusion in rotation. At the same time, our method demonstrates optimal stability and robustness in focusing on overlapping regions and gives correct decisions with the highest accuracy.

### v3

@@ Abstract
Currently, portable electronic devices are becoming more and more popular. For lightweight considerations, their fingerprint recognition modules usually use limited-size sensors. However, partial fingerprints have few matchable features, especially when there are differences in finger pressing posture or image quality, which makes partial fingerprint verification challenging. Most existing methods regard fingerprint position rectification and identity verification as independent tasks, ignoring the coupling relationship between them — relative pose estimation typically relies on paired features as anchors, and authentication accuracy tends to improve with more precise pose alignment. In this paper, we propose a novel framework for joint identity verification and pose alignment of partial fingerprint pairs, aiming to leverage their inherent correlation to improve each other. To achieve this, we present a multi-task CNN (Convolutional Neural Network)-Transformer hybrid network, and design a pre-training task to enhance the feature extraction capability. Experiments on multiple public datasets (NIST SD14, FVC2002 DB1_A & DB3_A, FVC2004 DB1_A & DB2_A, FVC2006 DB1_A) and an in-house dataset demonstrate that our method achieves state-of-the-art performance in both partial fingerprint verification and relative pose estimation, while being more efficient than previous methods.
@@ I Introduction
Figure 1: Partial fingerprint matching faces three challenges: sparse minutiae distribution, similar local texture and significant modal difference. This figure shows corresponding representative examples: (a) Three weakly-textured fingerprints that lack appropriate landmark feature points. (b) Local ridge patterns of fingerprints from different fingers have high similarity. (c) Fingerprint features in different skin conditions (in this case from left to right are dry, normal and wet respectively) may be missed (green) or incorrectly extracted (red). Visualized feature points are minutiae extracted by VeriFinger [9].
A natural idea for partial fingerprint matching is to transfer existing solutions for rolled or plain fingerprint recognition. Nevertheless, conventional algorithms usually rely on a sufficient number of key points to establish and compare spatial structural relationships [10, 11]. While certain studies use neural networks to extract fixed-length region descriptors [12, 13, 14, 15], consequently reducing the dependence on point features, these approaches require either large effective area or exceedingly precise alignment. Considering this, some researchers simultaneously input fingerprint images and extract interrelated features to fully utilize complete information [7, 16], which effectively improves the matching performance on partial fingerprints. However, to the best of our knowledge, almost all existing fingerprint matching methods (compared in Fig. 2) regard pose rectification and identity verification as independent tasks, ignoring the coupling relationship between them. Specifically, the estimation of relative pose typically relies on paired regional features as anchors, and the accuracy of authentication generally increases with enhanced precision in pose alignment. This is similar to a pictorial jigsaw puzzle where joint decision is suggested about whether (are they neighboring image fragments? ) and how (what’s the reassemble solution? ) any two pieces fit together [17, 18], in order to better balance robustness and speed in the overall process, rather than being made individually.
@@ II Related Work
According to different feature representation forms, existing one-to-one fingerprint matching algorithms can be mainly divided into three categories: key points based, fixed-length region descriptor based, and pairwise interrelated feature based. In the first two types of algorithms highly generalized features are stored and compared here (in other words, each fingerprint undergoes feature extraction only once) to enable efficient searches in large-scaled galleries. On the other hand, pairwise interrelated feature based methods could use richer information (features need to be re-extracted for each match) and is more suitable for confirmation scenarios with high precision requirements and few gallery / reference fingerprints, such as unlocking a mobile phone. Given the multitude of fingerprint matching algorithms, this rough classification may not comprehensively encompass all works. Nonetheless, we still endeavor to summarize them to the best of our knowledge. Fig. 2 shows the corresponding schematic diagrams of these three types of matching algorithms and ours.
@@ II-A Key Points Based
Minutiae are the most popular features used for fingerprint matching [3]. Researchers define minutiae descriptors using the attributes of minutiae themselves [10] and auxiliary information (such as orientation [27], period map [28], ridge [11] etc.). Considering the lack of reliable minutiae in latent fingerprints, Cao et al.[29] uniformly sample images into small patches and used a network to extract fixed length vector to describe virtual minutiae. However, as explained in Fig. 1, these features are not discriminative enough in partial fingerprints. On the other hand, some studies introduce universal key points in computer vision field into fingerprints [30, 6]. Features extracted in this way are more densely distributed, but at the same time more sensitive to unreliable image details and therefore easily confused. There are also algorithms that exploit Level-3 features from high-resolution images [31, 32], which are relatively expensive to deploy in most commercial applications. After getting key point sets, correspondence between them will be established based on the similarity of features and spatial structures, and then relative poses and matching scores are calculated.
@@ II-B Fixed-length Region Representation Based
Early studies use hand-crafted statistics to characterize fingerprint regions and convert them into fixed length vectors [33, 34, 35]. This feature form significantly improves the speed of large database indexing at the expense of reduced accuracy, and is usually used for rough searching. Engelsma et al.[12] propose an end-to-end network to extract fixed-length global representations of minutiae and textures, which greatly improves the performance of fixed-length representation methods. Inspired by this, researchers have conducted more exploration and improvement of fixed-length representation methods based on deep learning [36, 14, 37, 15]. In addition, Gu et al.[13] conduct dense sampling on complete fingerprints and compare the fixed length descriptors of each patch, ultimately making a comprehensive decision. These algorithms usually require pose rectification in advance, such as adaptively adjusting images through a spatial transformation layer [12, 15], rectifying absolute pose of fingerprints through a separate network [37], or exploiting the positioning relationships implicit in minutiae [14]. The error of pose estimation is tolerated to a certain extent because the extracted features are a generalization of regional characteristics. However, current methods may struggle to address partial fingerprints, as the pose estimation on them is often unstable.
@@ II-C Pairwise Interrelated Feature Based
Compared with key points or fixed-length descriptors, gray-level features of pixels contain raw information about local textures. Traditional algorithms use image correlation for fingerprint verification [38, 39], which can be affected by skin distortion and has limited discriminating ability. Dense fingerprint rectification and registration can reduce the impact of skin distortion on image correlation [40, 41], but simple correlation cannot separate genuine and impostor matches very well. With the development of deep learning, researchers apply neural networks to directly extract and verify interrelated features from two input images, significantly improving the verification performance [42, 43, 44, 7]. Specifically, He et al.[7] use multi-rotation and multi-size cropped image pairs as combined inputs to further assist the network in understanding scenarios with large spatial transformations or small overlapping areas. It should be noted that a significant difference between this type of method and methods based on fixed-length region descriptor is that the relative pose of a certain fingerprint is rectified instead of the respective absolute pose. Features of such methods need to be compared with precise spatial correspondence and thus have stricter requirements for alignment. Obviously, the verification performance will significantly decline with inaccurate pose estimation [7].
@@ III Method
In this paper, we propose JIPNet, a CNN-Transformer hybrid network, to jointly estimate the authentication probability and relative pose of partial fingerprint pairs. Fig. 3 gives the complete flowchart of our proposed algorithm. The overall process can be divided into four phases, which are indicated by different color bars in schematic: independent feature extraction (red), interrelated feature extraction (gold), identity verification (blue), and pose prediction (green). For an input fingerprint pair, our algorithm first extracts the respective features in parallel using convolutional blocks with shared weights (in Stem and Stages 1, 2). These two sets of features are then concatenated and fed to transformer blocks in Stage 3 (along with some extra microprocessing) for sufficient information interaction both locally and globally. Finally, respective output layers utilize the aggregated information to predict classification probabilities for authentication (Stage 4 - C and Classifier) and regression values for relative poses (Stage 4 - R and Regressor). It should be noted that these two tasks are performed in an integrated manner instead of independently as previous methods because we hope to utilize the mutually beneficial coupling relationship between each other, inspired by works on jigsaw puzzles [17, 18]. In particular, pose-related regression layers could be pruned to improve efficiency since only identity identification is necessary in most applications.
@@ III-B Interrelated Feature Extraction
The hybrid strategy of using early CNN and subsequent transformer to aggregate local details and global information has achieved great success in multiple fields [19, 20, 21]. Inspired by these, we utilize transformers to capture long-range dependencies within and across paired features, thereby ensuring sufficient information interaction. The paired features $f_{1}$ and $f_{2}$ are first biased with a sinusoidal positional encoding term [19] as follows:
@@
$PE_{(x,y,4i+3)}=\cos\left(y/10000^{4i/d}\right)\;,$
where $x$ and $y$ denote the 2D position, $i$ and $d$ denote the current and total dimensions. Features with added absolute positional information are then spatially flattened into 1D sequences and fed into stacked transformers with interleaved self- and cross- attention. It should be noted that the next two attention blocks have the same structure, including residual-connected Multi-Head Attention (MHA) and Feed-Forward Network (FFN) [48], and only the data streams are differentiated by their respective expected effects. Specifically, let $\operatorname{SA}\left(f_{i},f_{j}\right)$ (Self-Attention) and $\operatorname{CA}\left(f_{i},f_{j}\right)$ (Cross-Attention) represent the corresponding sub-blocks in Fig. 3, the complete process of transformer block is defined as:
@@ III-C Identification Verification & Pose Prediction
Similar to the previous phase, features are separately fed into stacked identical transformer blocks with head of $12$ to prepare for the corresponding task head. We define identity verification as a binary classification task, that is, judging whether the input image pair comes from a same finger. Let $\mathcal{F}$, $\operatorname{BN}$, $\operatorname{Swish}$ and $\operatorname{AvgPool}$ denote fully connected layer, batch normalization, swish activation and adaptive average pooling respectively, the classification probability $p$ is calculated as:
@@ V-E Visual Analysis
In order to provide more specific and intuitive comparison with interpretability, we employ occlusion sensitivity [16] to approximate the contribution of distinct local areas to the overall matching judgment. Fig. 10 shows four representative visualization results. Query fingerprint in images pairs is centered and search fingerprint is rigidly aligned based on the ground truth as background. Subsequently, a $32\times 32$ mask is slid and applied to search fingerprints and the differences between scores before and after occlusion are calculated, which are normalized and overlaid as the heatmap of search fingerprint. Considering that there is no reference relationship between the matching scores directly obtained (or mapped) by different algorithms, we convert these scores into a comparable probability form and attached them in corresponding visualization results. Specifically, given the overall distributions (shown in Fig. 7), the verification accuracy can be expressed as:
$P\left(y_{1}|s\right)$ $=\frac{P\left(s|y_{1}\right)P\left(y_{1}\right)}{P\left(s|y_{1}\right)P\left(y_{1}\right)+P\left(s|y_{0}\right)P\left(y_{0}\right)}$ (15)
$=\frac{f_{\mathrm{1}}\left(s\right)P\left(y_{1}\right)}{f_{\mathrm{1}}\left(s\right)P\left(y_{1}\right)+f_{\mathrm{0}}\left(s\right)P\left(y_{0}\right)}\;,$
where $s$ is the genuine matching score, $P$ and $f$ represent the probability and probability density, $y$ represents the identity to be verified (1 is the same, 0 is different). It can be seen that algorithms based on key points [6, 9] and region representation [12, 13] expose their respective shortcomings when there is a lack of sufficient features in overlapping areas (row 1) and significant differences in ridge texture of non overlapping areas (row 2). PFVNet [7] has clear advantages in both cases, even with minimal overlap (row 3). However, it fails in the last example with confusion in rotation. At the same time, our method demonstrates optimal stability and robustness in focusing on overlapping regions and gives correct decisions with the highest accuracy.